# A poset game on submonoids of additively indecomposable ordinals

### v2

@@ 3. Chomp on $(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})$
Let $\sigma=\delta+1\geq 2$. For every set $\{p_{1},\dots,p_{n}\}$ of natural numbers with $\textnormal{min}\{p_{1},\dots,p_{n}\}=r$ and for every $n\in\mathcal{S}^{1}$, the set
$\mathcal{S}^{\sigma}_{n}:=\{\alpha\in\mathcal{S}^{\sigma}\mid\omega^{\delta}\cdot n\leq_{\textnormal{Ord}}\alpha<_{\textnormal{Ord}}\omega^{\delta}\cdot(n+r)\}$
with the order $\leq_{\mathcal{S}^{\sigma}}\upharpoonright\mathcal{S}^{\sigma}_{n}$ is isomorphic to $(\mathcal{S}^{\delta};\leq_{\mathcal{S}^{\delta}})$. In addition, $\alpha<_{\mathcal{S}^{\sigma}}\omega^{\delta}\cdot(n+r)$ for every $\alpha\in\mathcal{S}^{\sigma}_{n}$.
@@
By definition of $r$, there is no $a\in\omega$ such that $n<_{\mathcal{S}^{1}}a<_{\mathcal{S}^{1}}n+r$, then, every $\alpha\in\mathcal{S}^{\sigma}_{n}$ has the form $\omega^{\delta}\cdot n+\beta$, with $\beta<\omega^{\delta}$. In fact, the function $f:\mathcal{S}^{\delta}\to\mathcal{S}^{\sigma}_{n}$ given by $f(\beta)=\omega^{\delta}\cdot n+\beta$ is an isomorphism.
The last part is clear because $\omega^{\delta}\cdot(n+r)=(\omega^{\delta}\cdot n+\beta)+\omega^{\delta}\cdot r$ for all $\beta\in\mathcal{S}^{\delta}$.∎
With these two lemmas, we can get a good idea of the form of the posets generated by natural numbers. As we can see in the examples in Figure 1, for any successor ordinal $\delta+1$, the shape is the same as for $\mathcal{S}^{1}$, but replacing each point with the poset $\mathcal{S}^{\delta}$.
@@
Lemma 3.3.
For every ordinal $\sigma$, Chomp on $\mathcal{S}^{\sigma}=\langle p_{1},\dots,p_{n}\rangle^{\sigma}$ is a finite game. Therefore, there must be a winning strategy on every poset of that form.
Proof.
Since $\alpha\leq_{\mathcal{S^{\sigma}}}\beta$ implies $\alpha\in\beta$, there are no infinite descending chains. We now show by transfinite induction that there are no infinite antichains.
The case $\sigma=0$ is trivial and the case $\sigma=1$ is also known (see [GMK18, Remark 2.3]). If this is true for $\sigma$ but not for $\sigma+1$, then, by Lemma 3.2 we could find an infinite antichain of the form $\{\omega^{\sigma}\cdot n_{i}+\alpha\mid i\in\omega,\alpha\in\mathcal{S}^{\sigma}\}$ where $n_{i}\not=n_{j}$ when $i\not=j$. But by Lemma 3.1 this is impossible, since there would exist an infinite antichain in $\mathcal{S}^{1}$. For $\sigma$ a limit ordinal, an infinite antichain in $\mathcal{S}^{\sigma}$ must be contained in some $\mathcal{S}^{\alpha}\subset S^{\sigma}$ for some $\alpha\in\sigma$, which contradicts the induction hypothesis. ∎
This, however, fails in general once we move to more general sets of generators.
@@
generate the following antichains
@@
A set $\Gamma$ of ordinals will be called a well-partial order generator ($\mathsf{wpog}$ for short) if the poset structure on $\langle\Gamma\rangle^{\sigma}$ is a well-partial order for every $\sigma$ such that $\Gamma\subset\omega^{\sigma}$.
@@
From what we saw at the beginning, an antichain in $\langle\Gamma\rangle^{\sigma+1}$ would give us an antichain of the form $A=\{\alpha\otimes\omega^{\gamma}\oplus\xi_{j}\mid j\in\omega\}$ where $\Xi=\{\xi_{j}\mid j\in\omega\}$ is some subset of $\langle\Gamma\rangle^{\sigma}$ and $\alpha\in\langle\Gamma\rangle^{\overline{e}(\Gamma)+1}$ is fixed and satisfies $\overline{e}(\omega^{\gamma}\otimes\alpha)=\sigma$. Assume that $\beta=\overline{e}(\Xi)\in\{\overline{e}(\xi_{j})\mid j\in\omega\}$, that is, for all $j\in\omega$, the first unfixed term of $\alpha\otimes\omega^{\gamma}\oplus\xi_{j}$ is of the form $\omega^{\beta}\cdot n_{j}^{\beta}$ whenever $\alpha_{i-1}<\beta<\alpha_{i}$, or $\omega^{\beta}\cdot(n_{j}^{\beta}+a_{i})$ when $\beta=\alpha_{i}$. If there are infinitely many $n_{j}^{\beta}$, we can apply the claim and obtain $\{\alpha_{\beta}\otimes\omega^{\gamma}\oplus\xi_{j}\mid j\in\omega\}\subset\langle\Gamma\rangle^{\sigma}$. This cannot be an antichain, so we may compare two elements $\alpha_{\beta}\otimes\omega^{\gamma}\oplus\xi_{j}$ and $\alpha_{\beta}\otimes\omega^{\gamma}\oplus\xi_{j^{\prime}}$. It is not hard to see that, consequently, we can also compare $\alpha\otimes\omega^{\gamma}\oplus\xi_{j}$ and $\alpha\otimes\omega^{\gamma}\oplus\xi_{j^{\prime}}$, which is absurd. Therefore, we may assume that there are only finitely many distinct $n_{j}^{\beta}$’s and, in particular, we may fix one of them and obtain a new infinite antichain $A^{\prime}\subseteq A$ whose first and second coefficients are fixed. If $\overline{e}(\Xi)>\overline{e}(\xi_{j})$ for all $j\in\omega$, we apply the second part of the claim in a similar way.
We may continue fixing coefficients, using the second part of the claim in the cases where $\alpha_{i}-\alpha_{i-1}\geq\omega$. After a finite number of steps, we will have fixed all coefficients between $\alpha_{1}$ and $\alpha_{m}$, and we will conclude that there is an infinite subset of $\{\xi_{j}\mid j\in\omega\}$ that forms an antichain, which is impossible. Hence, $A$ cannot be an antichain. Finally, if $\sigma$ is a limit ordinal, we proceed as in Lemma 3.3. ∎
It is worth noting that this theorem essentially shows that there exists a “Hanf number” for the property “the poset contains no antichains” and that, in fact, this number is the smallest possible ordinal. This, in part, motivates Question 4.9.
@@
Combining Theorem 3.7 and Lemma 3.3, we see that we only need to look for antichains between $\omega^{\alpha+1}$ and $\omega^{\alpha+2}$. Let $A\subset\langle\Gamma\rangle^{\alpha+2}$ be an infinite antichain; then, its elements can be written as
@@
This shows that the class of well-partial order generators is quite diverse.
4. Winning strategies
We start with an elementary but very general lemma that will be useful later.
Lemma 4.1.
Let $\Gamma$ be a $\mathsf{wpog}$. If player $A$ has a winning strategy on $(\mathcal{S}^{\alpha};\leq_{\mathcal{S}^{\alpha}})$, then $A$ also has a winning strategy on every poset $(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})$ with $\sigma>\alpha$. Hence, if player $B$ has a winning strategy on $(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})$, player $B$ also has a winning strategy on $(\mathcal{S}^{\alpha};\leq_{\mathcal{S}^{\alpha}})$ for every $\alpha<\sigma$.
If $\delta$ is a limit ordinal and $B$ has a winning strategy on $(\mathcal{S}^{\alpha};\leq_{\mathcal{S}^{\alpha}})$ for unboundedly many $\alpha$ less than $\delta$, then $B$ has a winning strategy on $(\mathcal{S}^{\delta};\leq_{\mathcal{S}^{\delta}})$.
Proof.
By Zermelo’s theorem, we only need to prove the first part. If the first move of $A$’s winning strategy on $(\mathcal{S}^{\alpha};\leq_{\mathcal{S}^{\alpha}})$ is to choose the ordinal $\beta$, then the winning move in $(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})$ is the same, because the poset remaining after this move is the same in both cases.
For the second part, note that, then, $B$ has a winning strategy on $(\mathcal{S}^{\alpha};\leq_{\mathcal{S}^{\alpha}})$ for every $\alpha<\delta$. Player $A$’s first move must necessarily be some ordinal $\beta<\omega^{\gamma}$ for some $\gamma<\delta$, so it will be a move in $\mathcal{S}^{\gamma}$, where $B$ has a winning strategy. ∎
Notation 4.2.
Given a $\mathsf{wpog}$ $\Gamma$, the previous lemma shows that if $B$ does not have a winning strategy on all the posets $(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})$, there must be an ordinal $\alpha$ such that $B$ has a winning strategy on $(\mathcal{S}^{\alpha};\leq_{\mathcal{S}^{\alpha}})$ but not on $(\mathcal{S}^{\gamma};\leq_{\mathcal{S}^{\gamma}})$ for all $\gamma\geq\alpha+1$. We let $\textnormal{ch}(\Gamma)$ denote $\alpha+1$.
In fact, with the following simple argument, we can talk about the “Hanf number” (see [Bal09, Ch. 4]) for the property “$B$ has a winning strategy”:
Proposition 4.3.
Fix an ordinal $\sigma$. There exists an ordinal $\xi$ such that, for every $\mathsf{wpog}$ $\Gamma\subset\omega^{\sigma}$, player $B$ has a winning strategy on $(\mathcal{S}^{\tau};\leq_{\mathcal{S}^{\tau}})$ for every ordinal $\tau\geq\sigma$ iff $B$ has a winning strategy on $(\mathcal{S}^{\xi};\leq_{\mathcal{S}^{\xi}})$.
Proof.
The idea is that, by fixing $\sigma$, we have fixed a set of classes of structures. Let us define
[…]
Take $\xi=\bigcup_{x\in S}\text{ch}(x)$ . Clearly, for every $x\subset\omega^{\sigma}$, if $B$ has a winning strategy on $\langle x\rangle^{\xi}$, then $x\not\in S$, that is, $B$ has a winning strategy on $\langle x\rangle^{\tau}$ for every $\tau\geq\sigma$. ∎
We can also prove that $\xi$ is the least possible for a specific family of generators, which we proceed to define.
@@ 4. Winning strategies
Unfortunately, Proposition 4.3 does not provide any information about $\xi$. Theorem 4.8 does give some information about this ordinal. But first, we need the following technical lemmas:
@@
For every finie set $\{\gamma_{1},\dots,\gamma_{n}\}=\Gamma\subset\textnormal{Ord}$, the correspoding class $\{(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})\mid\Gamma\subset\omega^{\omega^{\sigma}}\}$ is $\Sigma_{1}$-definable with $\gamma_{1},\dots,\gamma_{n}$ as parameters.
@@
Let $\Gamma$ be a finite $\mathsf{wpog}$. If $B$ does not have a winning strategy for every poset of the class $\{(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})\mid\Gamma\subset\omega^{\sigma}\}$ generated by $\Gamma$, then $\textnormal{ch}(\Gamma)<\omega_{1}$ if $\Gamma\subset\omega$ and $\textnormal{ch}(\Gamma)<|\max\{\Gamma\}|^{+}$ otherwise.
@@
By Lemma 4.7, we can apply [Bag12, Theorem 4.2] to the class of structures $C=\{(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})\mid\Gamma\subset\omega^{\omega^{\sigma}}\}$. Let $\kappa$ be equal to $\omega_{1}$ if $\Gamma\subset\omega$, and equal to $|\max\{\Gamma\}|^{+}$ otherwise. Then, for every $(\mathcal{S}^{\beta};\leq_{\mathcal{S}^{\beta}})\in C$, there exists $(\mathcal{S}^{\alpha};\leq_{\mathcal{S}^{\alpha}})\in C$ with $\alpha<\kappa$, such that there is some elementary embedding $j:\mathcal{S}^{\alpha}\to\mathcal{S}^{\beta}$. Take, for example, the least $\xi<\kappa$ for which there exists one of these elementary embeddings from $\mathcal{S}^{\xi}$ to $\mathcal{S}^{\sigma}$ with $\sigma\geq\kappa$. Now, if $B$ has a winning strategy on $(\mathcal{S}^{\xi};\leq_{\mathcal{S}^{\xi}})$, since there is an elementary embedding that preserves the sentences $\psi_{2n}$, we have $(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})\vDash\Psi$. Thus, by Lemma 4.1, $B$ has a winning strategy on every $(\mathcal{S}^{\alpha};\leq_{\mathcal{S}^{\alpha}})$ with $\alpha<\kappa$. Now we can transfer $\Psi$ to all structures above $\sigma$ using the corresponding elementary embeddings. Then, if $B$ has a winning strategy on $(\mathcal{S}^{\xi};\leq_{\mathcal{S}^{\xi}})$, $B$ also has a winning strategy on all posets in $C$ and therefore, again by Lemma 4.1, on all posets in $\{(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})\mid\Gamma\subset\omega^{\sigma}\}$, which proves that $\textnormal{ch}(\Gamma)<\kappa$. ∎
Although this theorem is quite general (and, in fact, its proof can be adapted for submonoids with the usual sum of ordinals instead of the natural sum), it is clear that it is far from optimal, and we believe that the bound provided can be significantly improved: very little information about the structure of the posets, which could be valuable, has been used so far. Moreover, here we have focused only on posets generated by finite sets.
Question 4.9.
@@
If the answer to Question 4.9 is affirmative, [GMK18, Theorem 6.5] would show us that the problem of determining the result of playing Chomp on any $(\mathcal{S}^{\sigma},\leq_{\mathcal{S}^{\sigma}})$ generated by natural numbers is decidable. However, the problem of finding explicit winning strategies in each case may still remain unresolved, leaving many open questions regarding the chomp game on these structures.

### v3

@@ 3.2. Algebraic approach
Let $\Gamma\not\subset\omega$ be a set of ordinals, and define
$\varepsilon:=\begin{cases}\left(\bigcup_{\gamma\in\Gamma}\overline{e}(\overline{e}(\gamma))\right)+1,&\text{if }\bigcup_{\gamma\in\Gamma}\overline{e}(\overline{e}(\gamma))\in\{\overline{e}(\overline{e}(\gamma))\mid\gamma\in\Gamma\},\\ \bigcup_{\gamma\in\Gamma}\overline{e}(\overline{e}(\gamma)),&\text{otherwise}.\end{cases}$
Then, every ordinal of the form $\omega^{\alpha}$, with $\alpha=\omega^{\epsilon_{1}}e_{1}+\cdots+\omega^{\epsilon_{n}}e_{n}+\omega^{\varepsilon}e$ (in Cantor normal form), is closed under $\otimes$ and $\oplus$ with elements of $\Gamma$; that is,
$\alpha_{1}\otimes\gamma_{1}\oplus\cdots\oplus\alpha_{k}\otimes\gamma_{k}\in\omega^{\alpha}$
for arbitrary $\alpha_{1},\dots,\alpha_{k}\in\omega^{\alpha}$ and $\gamma_{1},\dots,\gamma_{k}\in\Gamma$. Ordinals of this form will be called $\Gamma$-closed. Moreover, if $\omega^{\alpha}$ is $\Gamma$-closed, then every element $\beta\in\langle\Gamma\rangle^{\alpha}$ can be written as
$\beta=\omega^{\alpha_{1}}\otimes\beta_{1}+\cdots+\omega^{\alpha_{m}}\otimes\beta_{m}+\beta_{m+1},$
where the $\alpha_{i}$’s are $\Gamma$-closed, $\alpha>\alpha_{1}>\cdots>\alpha_{m}$, and $\beta_{1},\dots,\beta_{m+1}\in\langle\Gamma\rangle^{\omega^{\varepsilon}}$.
If $\Gamma\subset\omega$, any ordinal of the form $\omega^{\alpha}$ is $\Gamma$-closed and the corresponding decomposition was written in (1).
@@
The first part follows from Lemma 2.6; in particular, note that $\omega^{\omega^{\varepsilon}}$ is the least ordinal closed under $\oplus$ and $\otimes$ that contains $\Gamma$.
For the second part, we may factor out in $\beta$ common terms of the form $\omega^{\alpha_{i}}$ that are $\Gamma$-closed and chosen to be as large as possible. In this way, the coefficients multiplying each $\omega^{\alpha_{i}}$ are elements of $\langle\Gamma\rangle^{\omega^{\varepsilon}}$. Finally, we may replace $\oplus$ by ++, since in Cantor normal form the largest term of $\omega^{\alpha_{i+1}}\otimes\beta_{i+1}$ has an exponent smaller than the exponent of the smallest term of $\omega^{\alpha_{i}}\otimes\beta_{i}$.
And the case $\Gamma\subset\omega$ had already been considered. ∎
Proposition 3.9.
Let $\Gamma$ be a set of ordinals, let $\varepsilon$ be defined as in the previous lemma, and denote $(P;\leq_{P})=(\mathcal{S}^{\omega^{\varepsilon}};\leq_{\mathcal{S}^{\omega^{\varepsilon}}}).$ Then, for every poset $(P^{\sigma};\leq_{P^{\sigma}})$ there exists a $\Gamma$-closed ordinal $\omega^{\sigma^{\prime}}$ such that $(P^{\sigma};\leq_{P^{\sigma}})\cong(\mathcal{S}^{\sigma^{\prime}};\leq_{\mathcal{S}^{\sigma^{\prime}}}).$
Proof.
[…]
Nevertheless, observe that, in this context, nothing prevents us from considering a broader class of posets. Instead of studying the class of powers of a poset given by Definition 3.3, using an appropriate set $\Gamma$ of ordinals (see Definition 3.12), we may work with the corresponding class $\{(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})\mid\Gamma\subset\omega^{\sigma}\},$ which is more general. It is straightforward to verify that Lemma 3.5 also holds for this class of posets.
Notation 3.10.
Given an appropriate set $\Gamma$ of ordinals, if $B$ has not the winning strategy on all the posets $(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})$, we let $\textnormal{ch}(\Gamma)$ denote the ordinal $\xi+1$ such that player $B$ has the winning strategy on $(\mathcal{S}^{\xi};\leq_{\mathcal{S}^{\xi}})$ but not on $(\mathcal{S}^{\gamma};\leq_{\mathcal{S}^{\gamma}})$ for all $\gamma\geq\xi+1$.
3.3. Well partial orders
In this section we address a more practical question: the characterization of those sets of ordinals that give rise to well partial orders. Arguments such as those presented in Proposition 3.4 show that, if $\Gamma\subset\omega$, then any poset $(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})$ is a well partial order and, therefore, the game of Chomp can be played on these posets.
@@ 3.3. Well partial orders
This, however, fails in general once we move to more general sets of generators:
@@
generate the following infinite antichains
@@
A set $\Gamma$ of ordinals will be called a well partial order generator ($\mathsf{wpog}$ for short) if the poset structure on $\langle\Gamma\rangle^{\sigma}$ is a well partial order for every $\sigma$ such that $\Gamma\subset\omega^{\sigma}$.
@@
From what we saw at the beginning, an antichain in $\langle\Gamma\rangle^{\sigma+1}$ would give us an antichain of the form $A=\{\alpha\otimes\omega^{\gamma}\oplus\xi_{j}\mid j\in\omega\}$ where $\Xi=\{\xi_{j}\mid j\in\omega\}$ is some subset of $\langle\Gamma\rangle^{\sigma}$ and $\alpha\in\langle\Gamma\rangle^{\overline{e}(\Gamma)+1}$ is fixed and satisfies $\overline{e}(\omega^{\gamma}\otimes\alpha)=\sigma$. Assume that $\beta=\overline{e}(\Xi)\in\{\overline{e}(\xi_{j})\mid j\in\omega\}$; that is, for all $j\in\omega$, the first unfixed term of $\alpha\otimes\omega^{\gamma}\oplus\xi_{j}$ is of the form $\omega^{\beta}\cdot n_{j}^{\beta}$ whenever $\alpha_{i-1}<\beta<\alpha_{i}$, or $\omega^{\beta}\cdot(n_{j}^{\beta}+a_{i})$ when $\beta=\alpha_{i}$. If there are infinitely many $n_{j}^{\beta}$, we can apply the claim and obtain $\{\alpha_{\beta}\otimes\omega^{\gamma}\oplus\xi_{j}\mid j\in\omega\}\subset\langle\Gamma\rangle^{\sigma}$. This cannot be an antichain, so we may compare two elements $\alpha_{\beta}\otimes\omega^{\gamma}\oplus\xi_{j}$ and $\alpha_{\beta}\otimes\omega^{\gamma}\oplus\xi_{j^{\prime}}$. It is not hard to see that, consequently, we can also compare $\alpha\otimes\omega^{\gamma}\oplus\xi_{j}$ and $\alpha\otimes\omega^{\gamma}\oplus\xi_{j^{\prime}}$, which is absurd. Therefore, we may assume that there are only finitely many distinct $n_{j}^{\beta}$’s and, in particular, we may fix one of them and obtain a new infinite antichain $A^{\prime}\subseteq A$ whose first and second coefficients are fixed. If $\overline{e}(\Xi)>\overline{e}(\xi_{j})$ for all $j\in\omega$, we apply the second part of the claim in a similar way.
We may continue fixing coefficients, using the second part of the claim in the cases where $\alpha_{i}-\alpha_{i-1}\geq\omega$. After a finite number of steps, we will have fixed all coefficients between $\alpha_{1}$ and $\alpha_{m}$, and we will conclude that there is an infinite subset of $\{\xi_{j}\mid j\in\omega\}$ that forms an antichain, which is impossible. Hence, $A$ cannot be an antichain. Finally, if $\sigma$ is a limit ordinal, we proceed as in Proposition 3.4. ∎
It is worth noting that this theorem essentially shows that there exists a “Hanf number” for the property “the poset contains no antichains” and that, in fact, this number is the smallest possible ordinal. This, in part, motivates Question 4.6.
@@
Combining Theorem 3.14 with the fact that $p_{1},\dots,p_{n}$ do not generate infinite antichains, we see that we only need to look for antichains between $\omega^{\alpha+1}$ and $\omega^{\alpha+2}$. Let $A\subset\langle\Gamma\rangle^{\alpha+2}$ be an infinite antichain; then, its elements can be written as
@@
This shows that the class of well partial order generators is quite diverse.
4. Bounds for the function $\textnormal{ch}(\Gamma)$
We begin by proving, constructively, that $\textnormal{ch}(\Gamma)=1$ for a special family of generators, which we now proceed to define.
@@ 4. Bounds for the function $\textnormal{ch}(\Gamma)$
For our main result, we will need the following two lemmas.
@@
For every finie set $\{\gamma_{1},\dots,\gamma_{n}\}=\Gamma\subset\textnormal{Ord}$, the correspoding class $\{(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})\mid\text{ $\omega^{\sigma}$ is $\Gamma$-closed}\}$ is $\Sigma_{1}$-definable with $\gamma_{1},\dots,\gamma_{n}$ as parameters.
@@
Let $\Gamma$ be a finite $\mathsf{wpog}$. If $B$ does not have a winning strategy for every poset of the class $\{(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})\mid\Gamma\subset\omega^{\sigma}\}$ generated by $\Gamma$, then $\textnormal{ch}(\Gamma)<\omega_{1}$ if $\Gamma\subset\omega$ and $\textnormal{ch}(\Gamma)<|\bigcup\Gamma|^{+}$ otherwise.
@@
By Lemma 4.4, we can apply [Bag12, Theorem 4.2] to the class of structures $C=\{(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})\mid\text{$\omega^{\sigma}$ is $\Gamma$-closed}\}$. Let $\kappa$ be equal to $\omega_{1}$ if $\Gamma\subset\omega$, and equal to $|\bigcup\Gamma|^{+}$ otherwise. Then, for every $(\mathcal{S}^{\beta};\leq_{\mathcal{S}^{\beta}})\in C$, there exists $(\mathcal{S}^{\alpha};\leq_{\mathcal{S}^{\alpha}})\in C$ with $\alpha<\kappa$, such that there is some elementary embedding $j:\mathcal{S}^{\alpha}\to\mathcal{S}^{\beta}$. Take, for example, the least $\xi<\kappa$ for which there exists one of these elementary embeddings from $\mathcal{S}^{\xi}$ to $\mathcal{S}^{\sigma}$ with $\sigma\geq\kappa$. Now, if $B$ has a winning strategy on $(\mathcal{S}^{\xi};\leq_{\mathcal{S}^{\xi}})$, since there is an elementary embedding that preserves the sentences $\psi_{2n}$, we have $(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})\vDash\Psi$. Thus, by Lemma 3.5, $B$ has a winning strategy on every $(\mathcal{S}^{\alpha};\leq_{\mathcal{S}^{\alpha}})$ with $\alpha<\kappa$. Now we can transfer $\Psi$ to all structures above $\sigma$ using the corresponding elementary embeddings. Then, if $B$ has a winning strategy on $(\mathcal{S}^{\xi};\leq_{\mathcal{S}^{\xi}})$, $B$ also has a winning strategy on all posets in $C$ and therefore, again by Lemma 3.5, on all posets in $\{(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})\mid\Gamma\subset\omega^{\sigma}\}$, which proves that $\textnormal{ch}(\Gamma)<\kappa$. ∎
Although this theorem is quite general (and, in fact, its proof can be adapted for submonoids with the usual sum of ordinals instead of the natural sum), it is clear that it is far from optimal, and we believe that the bound provided can be significantly improved: very little information about the structure of the posets, which could be valuable, has been used so far. And, for families of posets with a more complex definition ($\Sigma_{n}$-definable for some $n\geq 2$), we would require large cardinals in the proof, which is even less optimal. Moreover, here we have focused only on posets generated by finite sets.
Question 4.6.
@@
If the answer to Question 4.6 is affirmative, [GMK18, Theorem 6.5] would show us that the problem of determining the result of playing Chomp on any $(\mathcal{S}^{\sigma},\leq_{\mathcal{S}^{\sigma}})$ generated by natural numbers is decidable. However, the problem of finding explicit winning strategies in each case may still remain unresolved, leaving many open questions regarding the chomp game on these structures.